OpenStudy (well):

Find Y,X. We have (2i -3)y + (2 +3i)x^2 = 2x - 3y + 15i. So we get : -3y + 2x^2 = 2x - 3y 2y + 3x^2 = 15 Next step will be probably modify on : 2x^2 - 2x = 0 3x^2 + 2y = 15 And that is my end. Dont know how to count with x^2 :/

OpenStudy (phi):

You are very far along. The first equation can be factored to find x

OpenStudy (anonymous):

2x^2 - 2x = 0 factor the x 2x(x-1) =0 x = 0;x = 1 sub it into the second equation to find y 3(0)^2 +2y = 15 y = 15/2 and 3(1)^2 + 2y = 15 y = 6 so you get values (0,15/2) and (1,6)

OpenStudy (well):

Thanks a lot. It takes me a while to realize. Thanks guys !